OpenStudy (anonymous):

the second side of a triangle is 5 inches more than the first side. The third side is 12 inches less than the first side. the perimeter of the triangle is 41 inches. Find each of the triangle.

OpenStudy (anonymous):

let x = first side x + 5 = second side x - 12 = third side x + x+5 + x - 12 = 41 3x -7 = 41 3x = 41 + 7 3X = 48 X = 48/3 X= 16 FIRST SIDE = 16 SECOND SIDE = 21 THIRDSIDE = 4

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

NO PROB ^^

OpenStudy (anonymous):

^^ you did this one already :)